# Towards exponentially-convergent simulations of extreme-mass-ratio inspirals: A time-domain solver for the scalar Teukolsky equation with singular source terms

### v2

@@ I Introduction
Black hole perturbation theory is a standard framework for studying a diverse range of gravitational phenomena, such as gravitational waves, quasi-normal modes, late-time Price tails, self-force effects, and linear stability of black hole solutions. The theory of such perturbations started with pioneering investigations by Regge and Wheeler Regge and Wheeler (1957); Zerilli (1970); Moncrief (1974) for Schwarzschild (nonspinning) black holes. The theory was later extended by Teukolsky to handle perturbations of the Kerr metric Teukolsky (1973). Small perturbations of a field with spin-weight $s$ evolve on the Kerr geometry according to the Teukolsky master equation (1).
@@
Despite the progress on time-domain methods for the source-free Teukolsky equation, particle-driven perturbations of the Kerr geometry remain challenging. To solve the distributionally-sourced Eq. (1), various regularized numerical approaches Engquist et al. (2005); Tornberg and Engquist (2004); Sundararajan et al. (2007, 2008) have been proposed. To our knowledge, all current treatments approximate the Dirac delta as a narrow Gaussian or discrete representations over an extended range of grid points Sundararajan et al. (2007, 2008); Harms et al. (2014). In both cases, these methods introduce small length scales (e.g., when the approximation is a narrow Gaussian), which can be a source of systematic error and will impose smaller timestep restrictions when evolved with an explicit time-stepping method. We note that an alternative effective-source approach can also be applied to this problem by allowing for analytic modifications to the source term Thornburg and Wardell (2017); Diener et al. (2012); Warburton and Wardell (2014).
@@
In this paper, we describe a discontinuous Galerkin (DG) method Reed and Hill (1973); Hesthaven and Warburton (2008); Cockburn et al. (2000); Cockburn (2001); Cockburn and Shu (1998) for solving the distributionally-sourced, $s=0$ Teukolsky equation describing scalar waves on Kerr. The main contribution of our work is to extend the numerical flux construction of Refs. Fan et al. (2008); Field et al. (2009, 2022) to the coupled $1+1$ system, Eq.  (II.4). Our DG method discretizes this coupled system written in fully first-order form and achieves global spectral accuracy everywhere in the computational domain including at the particle’s location. Furthermore, at future null infinity (where the far-field signal is recorded) our method is super-convergent, thereby allowing for extremely accurate waveforms for a comparatively lower-resolution numerical grid. As a by-product of our work, we also consider two systems that arise from different choices in the first-order reduction variables finding that certain choices are numerically problematic.
This paper is organized as follows. Section II derives the $1+1$ system in fully first-order form and discusses its hyperbolicity. The outer boundary, where we need to supply radiation outgoing boundary conditions and extract the far-field waveform, is handled using the method of hyperboloidal layers. The nodal discontinuous Galerkin method, and its generalization for incorporating Dirac delta distributions, are summarized in Sec. III. Section IV documents several experiments testing our method, including convergence tests against exact solutions in special cases where they are known, energy luminosities for circular orbits, and late-time tail behavior of the scalar field. Several appendices collect technical results and report on an alternative (and numerically problematic) formulation of the system. Appendix C shows our first-order system is symmetric hyperbolic.
@@ II.1 Teukolsky equation as a coupled 1+1D system
$-\left[\frac{(r^{2}+a^{2})^{2}}{\Delta}-a^{2}\sin^{2}\theta\right]\partial_{tt}\Psi-\frac{4Mar}{\Delta}\partial_{t\phi}\Psi$ (1)
@@
$-\left(s^{2}\cot^{2}\theta-s\right)\Psi=-4\pi\left(r^{2}+a^{2}\cos^{2}\theta\right)T,$
@@
Most time-domain numerical solvers do not directly discretize Eq. (1). Instead, due to the axisymmetry of the Kerr spacetime, a set of decoupled 2+1D equations can be derived by separating out the field’s azimuthal dependence. However, when the source term describes a point particle, both the original 3+1D and the 2+1D systems are numerically challenging. For example, for the 3+1D system, if $T\propto\delta(r-r_{p})\delta(\theta-\theta_{p})\delta(\phi-\phi_{p})$ then the solution is singular near $(r_{p},\theta_{p},\phi_{p})$. Moreover, the methods of Fan et al. (2008); Field et al. (2009), which exactly models the delta distribution as a modification to the numerical flux, are no longer directly applicable in 3+1D.
To overcome this issue, and to allow for exact treatment of the Dirac distribution, we follow Refs. Stein (2012); Dolan (2013); Barack and Giudice (2017) and derive a 1+1D coupled system of equations obeyed by the scalar field. We first expand the solution
@@
in terms of the ordinary scalar spherical harmonics $Y_{\ell m}(\theta,\phi)$. We follow the same definition and conventions as Ref. Thorne (1980) and, in particular, the harmonics are orthonormal when integrated over the sphere. Substituting Eq. (2) into Eq. (1) and using well-known properties of spherical harmonics we arrive at
@@ II.3 Reduction to first-order form
In these expressions and below we will use
@@
for convenience. If $\psi_{\ell m}$ solves the first-order system (II.3) it also solves the original second-order equation (II.1) provided the constraint $\phi_{\ell m}=\partial_{r_{*}}\psi_{\ell m}$. We note that alternative reduction variables could also be used. One such seemingly-reasonable choice ($\pi_{\ell m}=-\partial_{\tau}\psi_{\ell m}$ and $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$) is considered in App. D, and in Sec. IV we demonstrate that this choice appears to be numerically problematic. While we have not shown this system to be theoretically problematic, under certain conditions our preferred choice of reduction variables (15) leads to a symmetric hyperbolic (and hence well-posed) initial-boundary-value problem.
@@ II.4 Hyperboloidal layers
In hyperboloidal coordinates, Eq. (II.3) becomes  [^1] When enacting this coordinate transformation we do not let $\pi_{\ell m}$ and $\phi_{\ell m}$ transform. If we did, the resulting system would be the problematic one considered in App. D.
@@
where $\omega=\Omega^{2}/(\Omega-\rho\Omega^{\prime})$. To the left of the layer, where $\rho<R$, we have $H=0$ and so both Eq. (II.3) and Eq. (II.4) are identical. One important fact of this coordinate transformation, which we will use later on, is that the outgoing characteristics obey $\tau-\rho=t-r_{*}$. While the coordinate transformation is singular at $\rho=s$, the coefficients of each term in Eq. (II.4) are finite. This can be seen by noting that as $\rho\rightarrow s$ we have $\left(1-H\right)\sim\Omega^{2}\sim r_{*}^{-2}\sim r^{-2}$  Zenginoğlu (2011); Zenginoğlu and Khanna (2011).
@@ II.6 Matrix form
We now turn to writing Eq. (II.4) in matrix form, which will prove useful later on. Let $\ell_{\rm max}$ be the largest value of $\ell$ we will consider in the computation, that is we make an approximation $\psi_{\ell m}=0$ for $\ell>\ell_{\rm max}$. Noting that (i) the equation for odd and even $\ell$ modes decouple and (ii) each $m$ mode decouples, let us define a solution vector for the $-\ell$ mode, $U_{\ell_{\rm max}m}^{\rm even}=\left[\vec{\pi}_{\ell_{\rm max}m}^{\rm even},\vec{\phi}_{\ell_{\rm max}m}^{\rm even}\right]$, where
@@
The coupled system of equations (II.4)(b,c) take the form,
@@
where the components of $E$, $\hat{A}$, $\hat{B}$, and $\hat{V}$ can be read off from Eq. (II.4). The vector $\hat{G}(t)$ contains the coefficients of the distributional source term shown in Eq. (14). In this expression, we have dropped the $m$, $\ell_{\rm max}$, and $\{{\rm odd},{\rm even}\}$ labels from both the matrices and solution vector for clarity. The remaining set of equations, $\dot{\psi}_{\ell m}=-\pi_{\ell m}$, are trivially evolved along with the Eq. (25). Clearly, the size of these matrices change with $\ell_{\rm max}$ and $m$. For example, choosing $({\rm type},\ell_{\rm max},m)=({\rm even},8,0)$ will result in a $10$-by-$10$ system, while $({\rm type},\ell_{\rm max},m)=({\rm even},8,4)$ will yield a $6$-by-$6$ system. Inverting $E$, Eq. (25) becomes
@@ II.7 System hyperbolicity
We now consider the system’s hyperbolicity, which has important consequences for the numerics and is directly used in the construction of the numerical flux. As shown in App. B, the system (25) is (positive semi-definite) symmetric hyperbolic. If we had instead used $\pi_{\ell m}=-\partial_{\tau}\psi_{\ell m}$ and $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$ as our first-order reduction variables, the resulting system (D) is not symmetric hyperbolic even when $a=0$. In principle, this isn’t necessarily an issue, but we have found through extensive experimentation that our numerical scheme when applied to this alternative system shows a catastrophic loss of accuracy near $\rho=s$ (cf. Fig. 2 and Table 2). This alternative system and its hyperbolicity are discussed in App. D.
@@ II.8 Boundary conditions
Due to the hyperboloidal coordinate transformation, the wave speeds of the system, Eq.  (II.4) at the right physical boundary $\rho=s$ are either $1$ or 00. This means $\rho=s$ is an outflow boundary and hence no boundary conditions are needed. For the left boundary, we note that the system includes a mixture of incoming and outgoing characteristics and, furthermore, the potential is non-zero at the horizon when $m\neq 0$ and $a\neq 0$. We impose Sommerfeld boundary conditions setting the incoming characteristics to zero. This will lead to spurious reflection that can be mitigated by placing the left boundary at a sufficiently negative value of $\rho$ such that the left boundary is causally disconnected in the region of the computational domain that we care about. We note that this boundary issue can be partly solved by enacting so-called azimuthal transformations Krivan et al. (1996, 1997b), which we do not consider here except briefly in Sec. IV.3.2.
@@ II.9 Energy flux
One important application of the distributionally-sourced scalar wave equation is computing energy luminosities. Here we provide a formula for this quantity in terms of our system variables.
@@ III.1 The source-free method for $G=0$
We partition the spatial domain into $K$ non-overlapping subdomains defined by the partition points $\rho_{0}<\rho_{1}<\dots<\rho_{K}=s$ and denote $\mathsf{D}^{j}=[\rho_{j-1},\rho_{j}]$ as the $j^{\rm th}$ subdomain. In this one-dimensional setup, the points $\{\rho_{i}\}_{i=1}^{K-1}$ locate the internal subdomain interfaces, and we require one of them to be the location of the Dirac distribution. Since we are solving a one-dimensional problem, the subdomains are line segments and neighboring subdomains will intersect at a point.
@@ III.2 Inverting $\hat{E}$ and expressions at infinity
This symbolic approach works well for $\rho<s$. At $\rho=s$ the coordinate transformation (18) is singular yet, as remarked in Sec. II.4, the coefficients of the first-order differential equation (II.4) are well behaved. However, special care is needed at $\rho=s$. To better appreciate the issue at hand, consider setting $M=a=0$, then we end up with system (IV.1). The terms on the right-hand-side – these are $(1-H)$ and $V$ – behave like $r^{-2}$ as $r\rightarrow\infty$. Meanwhile, elements in the matrix $\hat{E}^{-1}$ behaves like $r^{2}$ as $r\rightarrow\infty$. Consequently, the coefficients and $\hat{E}^{-1}V$ need to be analytically supplied at $\rho=s$. This strategy continues to be applicable for the the general case (with non-zero values of $M$ and $a$) too, and some relevant expressions are provided in App. C.4. At grid points immediately to the left of $\rho=s$, the relevant coefficients in the partial differential operator might suffer from ill-conditioning (especially for $a\neq 0$ and large values of $\ell_{\rm max}$); we experimented with variable precision computations but found no benefit.
@@ IV Numerical experiments
The numerical experiments documented in Sec .IV.2 and IV.3 set $M=1$.
@@ IV.1 Verification of hyperboloidal layers: convergence and superconvergence
Our first experiment demonstrates the exponential convergence of our solver throughout the hyperboloidal layer. Upon setting $a=M=q=0$  [^2] Clearly this setup does not correspond to an EMRI model. However, the resulting system (IV.1) is an important special case for which exact solutions can be computed. This allows for unambiguous code tests that are impossible otherwise., the potential (10) becomes $V_{\ell m}=-\ell(\ell+1)/r^{2}$ and $r_{*}=r$. And so Eq. (II.1) is just the ordinary flatspace wave equation:
@@
The first-order system in hyperboloidal coordinates is given from Eq. (II.4):
@@
We solve Eq. (IV.1) (with $\ell=2$) on $\rho\in[1,50]$, with the layer at $R=30$ and using $P=4$, set the final time $T=50$, and choose a timestep $\Delta t=10^{-5}$ sufficiently small enough such that the Runge-Kutta’s temporal error is well below the spatial discretization error. Upon setting $f_{0}=2$, $c=1$, and $u_{0}=-10$, we take the initial data from the exact solution evaluated at $\tau=0$, which is (numerically) zero at the left boundary and the start of the layer $\rho=R$. At the left physical boundary point we choose numerical fluxes that weakly enforce simple Sommerfeld (outgoing) boundary conditions, which are sufficient for our purposes as the solution never impinges upon this boundary. Because of the property of the hyperboloidal layers, no boundary conditions are needed at $\rho=50$.
@@
Finally, Fig. 2 demonstrates the spectral convergence of our method as applied to Eq. (IV.1) (our preferred first-order reduced system; denoted “System 1” in the legend) and Eq. (D) (denoted “System 2” in the legend). We observe the problematic “System 2” loses convergence much sooner than “System 1”, in particular well before round-off error is expected to show up. While a complete understanding of this issue is still lacking, some observations are reported in App. D. For these experiments, we use a fixed number of $K=128$ subdomains and plot the error (57) as a function of $N$ and for each $N$. We have chosen a $\Delta t=0.00001$ to ensure the temporal error is well below the spatial discretization error.
Figure 2: Spectral (exponential) convergence of the $\ell=2$ hyperboloidal layers experiment described in Sec. IV.1. Errors have been computed from $\psi_{\rm num}(\tau,\rho=50)$, which is the signal that reaches future null infinity $\mathscr{I^{+}}$. Convergence results are reported for system 1 (dashed-red line; our preferred first-order reduced system (IV.1) that is symmetric hyperbolic) and system 2 (solid-blue line; an alternative first-order system (D) that is discussed in App. D). For system 1, we see excellent convergence with $N$ until round-off error effects become noticeable around $10^{-12}$. We find that our DG scheme solving system 2 is unable to obtain the expected accuracy that should be achievable by double precision. A complete understanding of this issue is still lacking; see App. D.

### v3

# Toward exponentially-convergent simulations of extreme-mass-ratio inspirals: A time-domain solver for the scalar Teukolsky equation with singular source terms

@@ I Introduction
Black hole perturbation theory is a standard framework for studying a diverse range of gravitational phenomena, such as gravitational waves, quasi-normal modes, late-time Price tails, self-force effects, and linear stability of black hole solutions. The theory of such perturbations started with pioneering investigations by Regge and Wheeler Regge and Wheeler (1957) and later important extensions by Zerilli and Moncrief Zerilli (1970); Moncrief (1974) for Schwarzschild (nonspinning) black holes. The theory was later extended by Teukolsky to handle perturbations of the Kerr metric Teukolsky (1973). Small perturbations of a field with spin-weight $s$ evolve on the Kerr geometry according to the Teukolsky master equation (II.1).
@@
Despite the progress on time-domain methods for the source-free Teukolsky equation, particle-driven perturbations of the Kerr geometry remain challenging. To solve the distributionally-sourced Eq. (II.1), various regularized numerical approaches Engquist et al. (2005); Tornberg and Engquist (2004); Sundararajan et al. (2007, 2008) have been proposed. To our knowledge, all current treatments approximate the Dirac delta as a narrow Gaussian or discrete representations over an extended range of grid points Sundararajan et al. (2007, 2008); Harms et al. (2014). In both cases, these methods introduce small length scales (e.g., when the approximation is a narrow Gaussian), which can be a source of systematic error and will impose smaller timestep restrictions when evolved with an explicit time-stepping method. We note that an alternative effective-source approach can also be applied to this problem by allowing for analytic modifications to the source term Thornburg and Wardell (2017); Diener et al. (2012); Warburton and Wardell (2014).
@@
In this paper, we describe a discontinuous Galerkin (DG) method Reed and Hill (1973); Hesthaven and Warburton (2008); Cockburn et al. (2000); Cockburn (2001); Cockburn and Shu (1998) for solving the distributionally-sourced, $s=0$ Teukolsky equation describing scalar waves on Kerr. The main contribution of our work is to extend the numerical flux construction of Refs. Fan et al. (2008); Field et al. (2009, 2022) to the coupled $1+1$ system, Eq.  (20). Our DG method discretizes this coupled system written in fully first-order form and achieves global spectral accuracy everywhere in the computational domain including at the particle’s location. Furthermore, at future null infinity (where the far-field signal is recorded) our method is super-convergent, thereby allowing for extremely accurate waveforms for a comparatively lower-resolution numerical grid. As a by-product of our work, we also consider two systems that arise from different choices in the first-order reduction variables finding that certain choices are numerically problematic.
This paper is organized as follows. Section II derives the $1+1$ system in fully first-order form and discusses its hyperbolicity. The outer boundary, where we need to supply radiation outgoing boundary conditions and extract the far-field waveform, is handled using the method of hyperboloidal layers. The nodal discontinuous Galerkin method, and its generalization for incorporating Dirac delta distributions, are summarized in Sec. III. Section IV documents several experiments testing our method, including convergence tests against exact solutions in special cases where they are known, energy luminosities for circular orbits, and late-time tail behavior of the scalar field. Several appendices collect technical results and report on an alternative (and numerically problematic) formulation of the system. Appendix C shows our first-order system is symmetric hyperbolic and well-posed.
@@ II.1 Teukolsky equation as a coupled 1+1D system
$-\left[\frac{(r^{2}+a^{2})^{2}}{\Delta}-a^{2}\sin^{2}\theta\right]\partial_{tt}\Psi-\frac{4Mar}{\Delta}\partial_{t\phi}\Psi$
@@
$-\left(s^{2}\cot^{2}\theta-s\right)\Psi=-4\pi\left(r^{2}+a^{2}\cos^{2}\theta\right)T,$ (1)
@@
Most time-domain numerical solvers do not directly discretize Eq. (II.1). Instead, due to the axisymmetry of the Kerr spacetime, a set of decoupled 2+1D equations can be derived by separating out the field’s azimuthal dependence. However, when the source term describes a point particle, both the original 3+1D and the 2+1D systems are numerically challenging. For example, for the 3+1D system, if $T\propto\delta(r-r_{p})\delta(\theta-\theta_{p})\delta(\phi-\phi_{p})$ then the solution is singular near $(r_{p},\theta_{p},\phi_{p})$. Moreover, the methods of Fan et al. (2008); Field et al. (2009), which exactly models the delta distribution as a modification to the numerical flux, are no longer directly applicable in 3+1D.
To overcome this issue and to allow for exact treatment of the Dirac distribution, we follow Refs. Stein (2012); Dolan (2013); Barack and Giudice (2017) and derive a 1+1D coupled system of equations obeyed by the scalar field. We first expand the solution
@@
in terms of the ordinary scalar spherical harmonics $Y_{\ell m}(\theta,\phi)$. We follow the same definition and conventions as Ref. Thorne (1980) and, in particular, the harmonics are orthonormal when integrated over the sphere. Substituting Eq. (2) into Eq. (II.1) and using well-known properties of spherical harmonics we arrive at
@@ II.3 Reduction to first-order form
In these expressions and below, we will use
@@
for convenience. If $\psi_{\ell m}$ solves the first-order system (16) it also solves the original second-order equation (II.1) provided the constraint $\phi_{\ell m}=\partial_{r_{*}}\psi_{\ell m}$ is satisfied. We note that alternative reduction variables could also be used. One such seemingly-reasonable choice ($\pi_{\ell m}=-\partial_{\tau}\psi_{\ell m}$ and $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$ based on the hyperboloidal coordinates introduced in Sec. II.4) is considered in App. D, and in Sec. IV we demonstrate that this choice appears to be numerically problematic.
@@ II.4 Hyperboloidal layers
In hyperboloidal coordinates, Eq. (16) becomes  [^1] When enacting this coordinate transformation we do not let $\pi_{\ell m}$ and $\phi_{\ell m}$ transform. If we did, the resulting system would be the problematic one considered in App. D.
@@
where $\omega=\Omega^{2}/(\Omega-\rho\Omega^{\prime})$. To the left of the layer, where $\rho<R$, we have $H=0$ and so both Eq. (16) and Eq. (20) are identical. One important fact of this coordinate transformation, which we will use later on, is that the outgoing characteristics obey $\tau-\rho=t-r_{*}$. While the coordinate transformation is singular at $\rho=s$, the coefficients of each term in Eq. (20) are finite. This can be seen by noting that as $\rho\rightarrow s$ we have $\left(1-H\right)\sim\Omega^{2}\sim r_{*}^{-2}\sim r^{-2}$  Zenginoğlu (2011); Zenginoğlu and Khanna (2011).
@@ II.6 Matrix form
We now turn to writing Eq. (20) in matrix form, which will prove useful later on. Let $\ell_{\rm max}$ be the largest value of $\ell$ we will consider in the computation, that is we make an approximation $\psi_{\ell m}=0$ for $\ell>\ell_{\rm max}$. Noting that (i) the equation for odd and even $\ell$ modes decouple and (ii) each $m$ mode decouples, let us define a solution vector for the $-\ell$ mode, $U_{\ell_{\rm max}m}^{\rm even}=\left[\vec{\pi}_{\ell_{\rm max}m}^{\rm even},\vec{\phi}_{\ell_{\rm max}m}^{\rm even}\right]$, where
@@
The coupled system of equations (20)(b,c) take the form,
@@
where the components of $E$, $\hat{A}$, $\hat{B}$, and $\hat{V}$ can be read off from Eq. (20). The vector $\hat{G}(t)$ contains the coefficients of the distributional source term shown in Eq. (14). In this expression, we have dropped the $m$, $\ell_{\rm max}$, and $\{{\rm odd},{\rm even}\}$ labels from both the matrices and solution vector for clarity. The remaining set of equations, $\dot{\psi}_{\ell m}=-\pi_{\ell m}$, are trivially evolved along with the Eq. (25). Clearly, the size of these matrices change with $\ell_{\rm max}$ and $m$. For example, choosing $({\rm type},\ell_{\rm max},m)=({\rm even},8,0)$ will result in a $10$-by-$10$ system, while $({\rm type},\ell_{\rm max},m)=({\rm even},8,4)$ will yield a $6$-by-$6$ system. Inverting $E$, Eq. (25) becomes
@@ II.7 System hyperbolicity
We now consider the system’s hyperbolicity, which has important consequences for the numerics and is directly used in the construction of the numerical flux. As shown in App. B, the system (25) is (positive semi-definite) symmetric hyperbolic and well-posed. If we had instead used $\pi_{\ell m}=-\partial_{\tau}\psi_{\ell m}$ and $\phi_{\ell m}=\partial_{\rho}\psi_{\ell m}$ as our first-order reduction variables, the resulting system (71) is not symmetric hyperbolic (although it is strongly hyperbolic) even when $a=0$. In principle, this isn’t necessarily an issue, but we have found through extensive experimentation that our numerical scheme when applied to this alternative system shows a catastrophic loss of accuracy near $\rho=s$ (cf. Fig. 2 and Table 2). The observed non-convergent behavior may be due to the fact that the problem’s boundary matrix  [^2] In one spatial dimension, the right boundary matrix is simply $P=A(\rho=s)$. See Sec. 5 of Ref. Sarbach and Tiglio (2012)., is not invertible. This alternative system and these issues are discussed in App. D.
@@ II.8 Boundary conditions
Due to the hyperboloidal coordinate transformation, the wave speeds of the system, Eq.  (20) at the right physical boundary $\rho=s$ are either $1$ or 00. This means $\rho=s$ is an outflow boundary, and hence no boundary conditions are needed. For the left boundary, we note that the system includes a mixture of incoming and outgoing characteristics and, furthermore, the potential is non-zero at the horizon when $m\neq 0$ and $a\neq 0$. We impose Sommerfeld boundary conditions setting the incoming characteristics to zero. This will lead to spurious reflection that can be mitigated by placing the left boundary at a sufficiently negative value of $\rho$ such that the left boundary is causally disconnected in the region of the computational domain that we care about. We note that this boundary issue can be partly solved by enacting so-called azimuthal transformations Krivan et al. (1996, 1997b), which we do not consider here except briefly in Sec. IV.3.2.
@@ II.9 Energy flux
One important application of the distributionally-sourced scalar wave equation is computing energy luminosities. Here, we provide a formula for this quantity in terms of our system variables.
@@ III.1 The source-free method for $G=0$
We partition the spatial domain into $K$ non-overlapping subdomains defined by the partition points $\rho_{0}<\rho_{1}<\dots<\rho_{K}=s$ and denote $\mathsf{D}^{j}=[\rho_{j-1},\rho_{j}]$ as the $j^{\rm th}$ subdomain. In this one-dimensional setup, the points $\{\rho_{i}\}_{i=1}^{K-1}$ locate the internal subdomain interfaces, and we require one of them to be the location of the Dirac distribution. Since we are solving a one-dimensional problem, the subdomains are line segments and neighboring subdomains will intersect at a point. We note that non-ciruclar orbits with $r_{p}(t)$ can also be handled by using a time-dependent coordinate transformation, which has been worked out in simpler settings Field et al. (2009).
@@ III.2 Inverting $\hat{E}$ and expressions at infinity
This symbolic approach works well for $\rho<s$. At $\rho=s$ the coordinate transformation (18) is singular yet, as remarked in Sec. II.4, the coefficients of the first-order differential equation (20) are well behaved. However, special care is needed at $\rho=s$. To better appreciate the issue at hand, consider setting $M=a=0$, then we end up with system (44). The terms on the right-hand-side – these are $(1-H)$ and $V$ – behave like $r^{-2}$ as $r\rightarrow\infty$. Meanwhile, elements in the matrix $\hat{E}^{-1}$ behaves like $r^{2}$ as $r\rightarrow\infty$. Consequently, the coefficients and $\hat{E}^{-1}V$ need to be analytically supplied at $\rho=s$. This strategy continues to be applicable for the the general case (with non-zero values of $M$ and $a$) too, and some relevant expressions are provided in App. C.4. At grid points immediately to the left of $\rho=s$, the relevant coefficients in the partial differential operator might suffer from ill-conditioning (especially for $a\neq 0$ and large values of $\ell_{\rm max}$); we experimented with variable precision computations but found no benefit.
@@ IV Numerical experiments
In this section, we perform several numerical experiments to benchmark our setup. In Sec. IV.1, we demonstrate spectral convergence and superconvergence of the algorithm when specialized to flatspace wave equation. In this simplified setting we can write down an exact analytical solution and perform unambiguous convergence tests that are impossible to do in Kerr. In Sec. IV.2 we verify known tail behavior at a finite distance and at future null infinity by numerically computing the solution’s late-time power-law decay behavior and its dependence on $\ell$. Finally, in Sec. IV.3, we numerically compute energy fluxes for circular orbits in Schwarzschild and Kerr geometries. We consider a variety of $(\ell,m)$ modes, orbital parameters, black hole spins, and also consider the impact of hyperboloidal coordinate transformation parameters and initial data on the quality of our results. The numerical experiments documented in Sec. IV.2 and IV.3 set $M=1$.
@@ IV.1 Verification of hyperboloidal layers: convergence and superconvergence
Our first experiment demonstrates the exponential convergence of our solver throughout the hyperboloidal layer. Upon setting $a=M=q=0$  [^3] Clearly this setup does not correspond to an EMRI model. However, the resulting system (44) is an important special case for which exact solutions can be computed. This allows for unambiguous code tests that are impossible otherwise., the potential (10) becomes $V_{\ell m}=-\ell(\ell+1)/r^{2}$ and $r_{*}=r$. And so Eq. (II.1) is just the ordinary flatspace wave equation:
@@
The first-order system in hyperboloidal coordinates is given from Eq. (20):
@@
We solve Eq. (44) (with $\ell=2$) on $\rho\in[1,50]$, with the layer at $R=30$ and using $P=4$, set the final time $T=50$, and choose a timestep $\Delta t=10^{-5}$ sufficiently small enough such that the Runge-Kutta’s temporal error is well below the spatial discretization error. Upon setting $f_{0}=2$, $c=1$, and $u_{0}=-10$, we take the initial data from the exact solution evaluated at $\tau=0$, which is (numerically) zero at the left boundary and the start of the layer $\rho=R$. At the left physical boundary point we choose numerical fluxes that weakly enforce simple Sommerfeld (outgoing) boundary conditions, which are sufficient for our purposes as the solution never impinges upon this boundary. Because of the property of the hyperboloidal layers, no boundary conditions are needed at $\rho=50$.
@@
Finally, Fig. 2 demonstrates the spectral convergence of our method as applied to Eq. (44) (our preferred first-order reduced system; denoted “System 1” in the legend) and Eq. (71) (denoted “System 2” in the legend). We observe the problematic “System 2” loses convergence much sooner than “System 1”, in particular well before round-off error is expected to show up. While a complete understanding of this issue is still lacking, some observations are reported in App. D. For these experiments, we use a fixed number of $K=128$ subdomains and plot the error (47) as a function of $N$ and for each $N$. We have chosen a $\Delta t=0.00001$ to ensure the temporal error is well below the spatial discretization error.
Figure 2: Spectral (exponential) convergence of the $\ell=2$ hyperboloidal layers experiment described in Sec. IV.1. Errors have been computed from $\psi_{\rm num}(\tau,\rho=50)$, which is the signal that reaches future null infinity $\mathscr{I^{+}}$. Convergence results are reported for system 1 (dashed-red line; our preferred first-order reduced system (44) that is symmetric hyperbolic) and system 2 (solid-blue line; an alternative first-order system (71) that is discussed in App. D). For system 1, we see excellent convergence with $N$ until round-off error effects become noticeable around $10^{-12}$. We find that our DG scheme solving system 2 is unable to obtain the expected accuracy that should be achievable by double precision. A complete understanding of this issue is still lacking; see App. D.